OpenStudy (anonymous):

dentity the graph of the equation 2x2 - 4xy + 2y2 - 7x + 7y = 0, then find , the angle of rotation.

OpenStudy (tkhunny):

The discriminant says \((-4)^{2} - 4(2)(2) = 16 - 16 = 0\) Shall we go with "Parabola"?

OpenStudy (anonymous):

yes then to find the angle I had \[\tan2\theta=-4/2-2\] but that doesn't make sense, what is the angle of rotation?

OpenStudy (tkhunny):

If you add the parentheses in the denominator, you get -4/0!!!! Gaaa!!! What could that mean? It means \(\theta = \pi/4\)

OpenStudy (anonymous):

which is 45 right?

OpenStudy (tkhunny):

Yes, 45º.

OpenStudy (anonymous):

thanks so much for the help!

OpenStudy (tkhunny):

Happy Conic Section Spinning!